A book is to be written by Britney Spears. Batman Books agrees to pay Britney $6 million for the rights to this not- yet- written memoir. According to one leading publisher, BatmanBooks could earn a profit of roughly $1.2 million if it sold 625,000 copies in hardcover. On the other hand, if it sold 375,000 copies, managers would lose about $1.3 million. Publishing executives stated that it was hard to sell more than 500,000 copies of a nonfiction hardcover book, and very exceptional to sell 1 million copies. Were Batman managers taking a substantial risk in publishing this book?


1
Expert's answer
2021-04-27T07:31:14-0400

In this case, the fixed cost is $6 million.


For the selling of 625,000 units:

Total contribution = Fixed cost + Profit

= 6 + 1.2 = $7.2 million


Contribution per unit = "\frac {\$7,200,000}{625,000} =\$11.52"


Break-even units = "\frac {Fixed cost} {Contribution per unit} = \frac {\$6,000,000} {\$11.52} = 520,833"


For the selling of 375,000 units:


Total contribution = Fixed cost "-"Loss

= 6"-"1.3 = $4.7 million


Contribution per unit = "\frac {\$4,700,000} {375,000} = \$12.53"


Break-even units ;"\frac {Fixed cost} {Contribution per unit} =\frac {\$6,000,000} {\$12.53} = 478,851"


Conclusion: No; there is no substantial risk of publishing.A publishing of more than 478,851 copies up to 500,000 copies could be sold out and can earn profit too; because as per the above calculation the break-even units is 478,851, anything above this would be profitable but it should not cross 500,000 because of the risk of non-selling. Therefore, there is still a chance of making the publication fruitful in normal way.